OpenStudy (anonymous):

In triangle ABC, AB = 7 and AC = 17. Find m

OpenStudy (anonymous):



OpenStudy (anonymous):

tan x = 17 / 7 x = arctan 17/7

OpenStudy (johnweldon1993):

Okay...well I'm assuming you know the trig functions....well here you will have to use the inverse trig functions sun^-1 , cos^-1 , and tan^-1 You use these when you have sides...but want to find angles.....just like you use sin, cos, and tan when you have an angle and a side and you want to find the other side... So...you have some angle *m*...a side opposite that angle....and a side adjacent to that angle..sounds like a tangent function...so this can be set up as \[\tan^{-1} \frac{ 17 }{ 7 } = ?\]

OpenStudy (anonymous):

67.6?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Thats what I get yeah

OpenStudy (anonymous):

ok thank you :)

OpenStudy (johnweldon1993):

anytime :)